40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 04:42:24 2011, kiwamu, Update, LSC, DRMI locked and some plans 
    Reply  Fri Oct 7 11:53:47 2011, Koji, Update, LSC, DRMI locked and some plans 
Message ID: 5627     Entry time: Fri Oct 7 04:42:24 2011     Reply to this: 5629
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: DRMI locked and some plans 

DRMI has been locked using the same RFPD selection as the old days (i.e. AS55_I, AS55_Q and REFL_I).(#4760)

But remember : this is just a beginning of several measurements and tests to characterize the central part.

 

Here is a list of the measurements and actions :

  - 3f locking related

      + Listing up the necessary RFPDs and their installations.

      + Calibration of the SRM actuator  => this is necessary to convert the sensing matrix into unit of [counts/m] or [W/m].

      + Measurement of the sensing matrix => check the performance of 3f signals. Also diagonalization of the LSC sensing matrix

      + Diagonalization of the output matrix.

      + Noise characterization of 3f PDs => confirm the noise are low enough to keep the lock of the central part

 - Power-recycling gain issue related (#5541)

     + Estimation of the mode matching efficiency => maybe we can use power-recycled ITMs to estimate it (?)

     + Implementation of auto alignment servos and scripts for MICH, PRCL and SRCL. => integrate it to the existing ASS model

     + Search for a possible loss factor

ELOG V3.1.3-